OpenStudy (anonymous):

what is the square root of -5

OpenStudy (anonymous):

and how did you find it

OpenStudy (anonymous):

This is an imaginary answer ...

OpenStudy (lollylau):

sqrt (5) i?

OpenStudy (anonymous):

since it is negative, take the negative sign out, since we do not know the square root of 5, we have \[\sqrt{5}-\]

OpenStudy (lollylau):

i = sqrt(-1) these are called imaginary numbers... written in the form of a+bi

OpenStudy (anonymous):

oppp\[-\sqrt{5} \]

OpenStudy (lollylau):

if drawn on a grid, the x-axis is the number line and the y-axis is the imaginary number line.

OpenStudy (anonymous):

it is 2.23i. i is the mathematical term for the square root of -1

OpenStudy (earthcitizen):

\[i \sqrt{5} =i2.2\]